the length and breadth of a garden are in the ratio of 3:2 if its length is 150 m then finds its area.
Answer by checkley79(3341)

L=150
3/2=150/X CROSS MULTIPLY.
3X=2*150
3X=300
X=300/3
X=100 ANS. FOR THE WIDTH.